OpenStudy (anonymous):

Translate the following situation into an equation. Do not solve. “On Monday afternoon Kelly withdrew 5% of the balance of her bank account to loan her friend some money. On Tuesday morning she deposited $100 into her bank account. She now has $25 more than she did on Monday morning. How much was in her bank account Monday morning?”

OpenStudy (anonymous):

x = amount on Monday morning x - .05x = amount on Monday afternoon 100 + x - .05x = amount on tuesday morning = twenty five more then monday morning = x +25 So..... 100 + x - .05x = x + 25

OpenStudy (anonymous):

omg! thank you. I have so much trouble with word problems. I appreciate the help. :-)

OpenStudy (anonymous):

sure, so did I for a long time. you really have to break them into parts.. with a lot of practice you'll get much better.